# Towards Automated Identification of Violation Symptoms of Architecture Erosion

### v3

@@ 7 Related Work
7.1 Code Review Comments
Code review comments contain massive knowledge related to software development, and a variety of studies analyze software defects and evolution through mining review comments and commit records. Han et al. [66] conducted an empirical study by manually checking 25,415 code review comments to investigate code smells identified in modern code review from four OSS projects. Their results show that most reviewers provide constructive suggestions to help developers fix code, and developers are willing to repair the smells through suggested refactoring operations. El Asri et al. [67] empirically investigated the impact of expressed sentiments on the code review duration and its outcome. They found that reviews with negative comments on average took longer time to complete than the reviews with positive/neutral comments. Kashiwa et al. [68] reported Self-Admitted Technical Debt (SATD) during code review by checking 156,372 review records from OpenStack and Qt. They found that 28-48% SATD are introduced during code review, and 20% SATD comments are created because of reviewers’ requests.
Besides, Uchôa et al. [69] investigated the impact of code review on the evolution of design degradation through mining and analyzing a plethora of code reviews from seven OSS projects. They found that there is a wide fluctuation of design degradation during the revisions of certain code reviews. Paixão et al. [70] explored how developers perform refactorings in code review, and they found that refactoring operations are most often used in code reviews that implement new features. Besides, they observed that the refactoring operations were rarely refined or undone along the code review, and such refactorings often contribute to new code smells and bugs.
@@ 7.1 Code Review Comments
7.2 Architecture Violations
Over the past decades, there have been various investigations on architecture violations. Brunet et al. [71] performed a longitudinal study to explore the evolution of architectural violations in 19 bi-weekly versions of four open source systems. They investigated the life cycle and location of architecture violations over time by comparing the intended and recovered architectures of a system. They found that architectural violations tend to intensify as software evolves and a few design entities are responsible for the majority of violations. More interestingly, some violations seem to be recurring after being eliminated. Mendoza et al. [13] proposed a tool called ArchVID based on model-driven engineering techniques for identifying architecture violations from source code; this tool supports recovering and visualizing the implemented architecture.
Besides, Terra et al. [72] reported their experience in fixing architectural violations. They proposed a recommendation system ArchFix that provides refactoring guidelines for developers and maintainers to repair architectural violations in the module architecture view of object-oriented systems. Their results show that their approach can provide correct recommendations for 75% and 79% of the detected architecture violations on two selected systems, respectively. Maffort et al. [73] proposed an approach to check architecture conformance for detecting architecture violations based on defined heuristics. They claimed that their approach relies on the defined heuristic rules and can rapidly raise architecture violation warnings.
@@ 7.2 Architecture Violations
7.3 Analyzing Software Repositories with Machine Learning and Deep Learning
The application of ML and DL techniques to extract informative knowledge from various artifacts during software development is gaining popularity in the software engineering community. In recent years, a plethora of empirical studies have been conducted to investigate automated techniques for supporting software engineering tasks. Khalajzadeh et al. [53] conducted an empirical study by extracting and manually analyzing app reviews from 12 OSS projects and developed ML and DL models to automatically detect and classify human-centric issues from app reviews and discussions. Their results show that automated techniques can help developers to recognize and appreciate human-centric issues of end-users more easily. Nasab et al. [52] developed 15 ML and DL models to automatically identify security discussions, and then they collected practitioners’ feedback through a validation survey and confirmed the promising applications of the models in practice.
AlOmar et al. [74] proposed an approach to automatically identify and classify self-admitted refactoring in commit messages. They compared their approach with pattern-based and simple random baselines, and the results show that a good performance and a relatively small training dataset is sufficient to classify self-admitted refactoring commits. Hey et al. [75] proposed a novel approach, named NoRBERT, for requirements classification by using the transfer learning capabilities of BERT. They applied their approach to different tasks in the domain of requirements classification, and the results show that their approach improves requirements classification and can be applied to unseen projects with convincing results.
In this work, word embeddings were utilized as word representations and as the input for training ML and DL-based classifiers, as prior studies have demonstrated the effectiveness of word embeddings in capturing rich semantic and syntactic features of words [61]. We employed ML/DL-related techniques to promote the automated identification of violation symptoms. Although much work has been done on automatic text identification, there is currently no automatic way to identify architecture violations in code review comments. Our approach is not restricted to specific textual information, so future studies can replicate and extend our work with other artifacts.
8 Conclusions and Future Work
@@ Acknowledgements
We would like to thank all the practitioners who participated in our survey and interviews. This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311.

### v4

@@ 7. Related Work
7.1. Code Review Comments
Code review comments contain massive knowledge related to software development, and a variety of studies analyze software defects and evolution through mining review comments and commit records. Han et al. (Han2022cmd, ) conducted an empirical study by manually checking 25,415 code review comments to investigate code smells identified in modern code review from four OSS projects. Their results show that most reviewers provide constructive suggestions to help developers fix code, and developers are willing to repair the smells through suggested refactoring operations. El Asri et al. (ElAsri2019ess, ) empirically investigated the impact of expressed sentiments on the code review duration and its outcome. They found that reviews with negative comments on average took longer time to complete than the reviews with positive/neutral comments. Kashiwa et al. (Kashiwa2022SATDmcr, ) reported Self-Admitted Technical Debt (SATD) during code review by checking 156,372 review records from OpenStack and Qt. They found that 28-48% SATD are introduced during code review, and 20% SATD comments are created because of reviewers’ requests.
Besides, Uchôa et al. (Uchoa2020hdm, ) investigated the impact of code review on the evolution of design degradation through mining and analyzing a plethora of code reviews from seven OSS projects. They found that there is a wide fluctuation of design degradation during the revisions of certain code reviews. Paixão et al. (Paixao2020bti, ) explored how developers perform refactorings in code review, and they found that refactoring operations are most often used in code reviews that implement new features. Besides, they observed that the refactoring operations were rarely refined or undone along the code review, and such refactorings often contribute to new code smells and bugs.
@@ 7.1. Code Review Comments
7.2. Architecture Violations
Over the past decades, there have been various investigations on architecture violations. Brunet et al. (Brunet2012av, ) performed a longitudinal study to explore the evolution of architectural violations in 19 bi-weekly versions of four open source systems. They investigated the life cycle and location of architecture violations over time by comparing the intended and recovered architectures of a system. They found that architectural violations tend to intensify as software evolves and a few design entities are responsible for the majority of violations. More interestingly, some violations seem to be recurring after being eliminated. Mendoza et al. (Mendoza2021avd, ) proposed a tool called ArchVID based on model-driven engineering techniques for identifying architecture violations from source code; this tool supports recovering and visualizing the implemented architecture.
Besides, Terra et al. (Terra2015rsr, ) reported their experience in fixing architectural violations. They proposed a recommendation system ArchFix that provides refactoring guidelines for developers and maintainers to repair architectural violations in the module architecture view of object-oriented systems. Their results show that their approach can provide correct recommendations for 75% and 79% of the detected architecture violations on two selected systems, respectively. Maffort et al. (Maffort2016mav, ) proposed an approach to check architecture conformance for detecting architecture violations based on defined heuristics. They claimed that their approach relies on the defined heuristic rules and can rapidly raise architecture violation warnings.
@@ 7.2. Architecture Violations
7.3. Analyzing Software Repositories with Machine Learning and Deep Learning
The application of ML and DL techniques to extract informative knowledge from various artifacts during software development is gaining popularity in the software engineering community. In recent years, a plethora of empirical studies have been conducted to investigate automated techniques for supporting software engineering tasks. Khalajzadeh et al. (Khalajzadeh2022sda, ) conducted an empirical study by extracting and manually analyzing app reviews from 12 OSS projects and developed ML and DL models to automatically detect and classify human-centric issues from app reviews and discussions. Their results show that automated techniques can help developers to recognize and appreciate human-centric issues of end-users more easily. Nasab et al. (Nasab2021ais, ) developed 15 ML and DL models to automatically identify security discussions, and then they collected practitioners’ feedback through a validation survey and confirmed the promising applications of the models in practice.
AlOmar et al. (AlOmar2021SAR, ) proposed an approach to automatically identify and classify self-admitted refactoring in commit messages. They compared their approach with pattern-based and simple random baselines, and the results show that a good performance and a relatively small training dataset is sufficient to classify self-admitted refactoring commits. Hey et al. (Hey2020norbert, ) proposed a novel approach, named NoRBERT, for requirements classification by using the transfer learning capabilities of BERT. They applied their approach to different tasks in the domain of requirements classification, and the results show that their approach improves requirements classification and can be applied to unseen projects with convincing results.
In this work, word embeddings were utilized as word representations and as the input for training ML and DL-based classifiers, as prior studies have demonstrated the effectiveness of word embeddings in capturing rich semantic and syntactic features of words (Ren2019nndSATD, ). We employed ML/DL-related techniques to promote the automated identification of violation symptoms. Although much work has been done on automatic text identification, there is currently no automatic way to identify architecture violations in code review comments. Our approach is not restricted to specific textual information, so future studies can replicate and extend our work with other artifacts.
8. Conclusions and Future Work
@@ Acknowledgements
We would like to thank all the practitioners who participated in our survey and interviews. This work has been partially supported by the National Natural Science Foundation of China (NSFC) with Grant No. 62172311 and 62402348.